OpenStudy (anonymous):

Can you help me what's 1/7 + 5/6

OpenStudy (anonymous):

fractions?

OpenStudy (anonymous):

what do you think it is

OpenStudy (anonymous):

41/42

OpenStudy (anonymous):

try to help them figure it out. they wont learn if you flat out give them the answer

OpenStudy (anonymous):

I times the denominator by the each other and then timed the numerator by the same number as I timesed the denominators by. I would've reduced after that except I couldn't reduce.

OpenStudy (anonymous):

y id I'm sorry if my response is hard to understand, my computers acting wacky and I can't backspace or draw, otherwise I would have drawn it out with the explanation.

OpenStudy (mokeira):

\[\frac{ 1 }{ 7 }+\frac{ 5 }{ 6 }\] First thing to do is to find the LCM of 7 and 6 so that you can have a common denominator LCM= 42 \[\frac{ }{ 42}\] To get the numerator, divide 42 by each respective denominator and multiply by its respective numerator. \[\frac{ 6+35 }{ 42 }\] I hope you have understoo